# A Comprehensive Survey on Self-Supervised Learning for Recommendation

### v2

@@ 4.5.3. Generative Learning in CroRec
  • Generative Learning Paradigm in CroRec begins by employing deep generative latent variable models (i.e., variational auto-encoders) to encode the latent space of user preferences. Subsequently, it transitions to the denoising diffusion model to facilitate knowledge transfer across domains. Salah et al. propose RA/SA-VAE (Salah et al., 2021) to simultaneously fit the target observations and align the hidden space encoded by VAE with the source latent space. They introduce both rigid alignment and soft alignment techniques with varying degrees of user preference alignment. Furthermore, VDEA (Liu et al., 2022c) extends the alignment of latent spaces using variational autoencoding at both local and global levels, allowing for the exploitation of domain-invariant features across different domains, including both overlapped and non-overlapped users. Recently, DiffCDR (Xuan, 2024) leveraged denoising probabilistic models (DPMs) to process noisy data and generate denoised results, effectively transferring data between domains.
  • Generation Target. RA/SA-VAE and VDEA both adopt the user rating reconstruction paradigm to optimize the variational model. In these models, the generation target is the user ratings for items, represented as vectors in the model. The generated ratings include ratings for uninteracted items, which are subsequently used for ranking and recommendation. On the other hand, DiffCDR focuses on generating user features in the target domain by reversing the diffusion process conditioned on the corresponding user’s information in the source domain. These generated user features in the target domain are then utilized for recommendation purposes.
@@ 4.5.4. Adversarial Learning in CroRec
  • Adversarial Learning Paradigm. Several methods employ adversarial learning for cross-domain recommendation. For instance, Su et al. (Su et al., 2022) use a discriminator to identify domain-specific features, with gradients propagating to optimize the feature generator. Similarly, RecGURU (Li et al., 2022d) feeds encoded user representations from both domains into a discriminator for classification, enabling differentiable adversarial learning. Moreover, ACLCDR (Hsu et al., 2023) utilizes Deep Q-Learning to design a generator that enhances the interaction matrix with fake interactions, optimizing it using reward signals from downstream tasks. In addition, DA-CDR (Zhang et al., 2021b) encodes domain features of users and items to fool the discriminator, ensuring effective domain knowledge transfer through adversarial learning using a gradient reversal layer (GRL) for differentiable training. Furthermore, DA-DAN (Guo et al., 2023b) incorporates adversarial domain adaptation into unsupervised non-overlapping CroRec, with the discriminator classifying the domain of encoded user features for overall differentiable training.
  • Discrimination Target. Su et al. (Su et al., 2022)’s domain adaptation method uses features from source and target domains to discriminate between them, forcing the generator to encode domain-invariant user preferences, thereby confusing the discriminator. RecGURU adopts a similar approach, discriminating encoded user representations from both domains to generalize user representation learning across domains. In ACLCDR, the augmented interaction matrix from DQN is used for inference in subsequent models, with the recommendation model acting as an implicit discriminator. DA-CDR employs a discriminator to classify domain features of users/items, encouraging the generator to encode additional domain-shared information for cross-domain recommendation. Lastly, DA-DAN encodes user-specific item sequence representations as domain features for the discriminator’s classification task, enabling the learning of domain-invariant features for recommendation.
@@ 4.6. Bundle Recommendation
GL Method Venue Year Generative Learning Paradigm Generation Target
DGMAE (Ren et al., 2023a) 2023 SIGIR Maksed Autoencoding Edge in User-Item Graph
@@
CubeRec (Chen et al., 2022d) 2022 SIGIR Model-based Score-based Explicit
@@ 4.6.2. Contrastive Learning in BunRec
In bundle recommendation with contrastive learning (Table 7), the relationship between items and bundles, along with user-item interactions, is utilized to form user-bundle relationships. Various methods are then designed to create views based on these relationships.
  • View Creation. MIDGN (Zhao et al., 2022) employs neural models to encode distinct representations for each user and bundle, treating them as separate views. It incorporates a graph disentangle module to encode intent-aware representations and utilizes a user-bundle graph to encode cross-view representations with LightGCN. Additionally, CrossCBR (Ma et al., 2022) initially builds the user-bundle graph by leveraging the user-item and bundle-item graphs. Subsequently, it encodes user and bundle representations using graph augmentation and embedding augmentation techniques in the bundle and item views. As a result, each user and bundle node possesses two views for contrastive learning, similar to the approach adopted in MIDGN.
  • Pair Sampling. Both MIDGN and CrossCBR encode two distinct views for each user and bundle in the dataset. As a result, the two views of a specific user/bundle node naturally form the positive pair, while other views from different nodes serve as negative samples for contrast comparison.
  • Contrastive Objective. MIDGN and CrossCBR both utilize the InfoNCE-based objective to optimize their models through self-supervised learning signals.
4.6.3. Generative Learning in BunRec
In BunRec, generative learning (Table 7) has adopted the concept of a graph-masked auto-encoder. The approach (i.e., DGMAE (Ren et al., 2023a)) is illustrated as follows.
  • Generative Learning Paradigm. In DGMAE, a teacher GNN model is first trained using user-bundle interactions, and then its knowledge is distilled into a student graph-masked auto-encoder model. This learning paradigm in DGMAE follows the principles of mask autoencoding.
  • Generation Target. DGMAE employs an adaptive masking strategy on the prominent edges within the user-item graph for the purpose of reconstruction. Specifically, edges with lower sparsity scores are prioritized for masking in DGMAE, increasing the reconstruction challenge and improving the model’s robustness against sparsity.
@@ 4.7.2. Contrastive Learning in GroRec
  • View Creation. GroupIM (Sankar et al., 2020) pioneers the use of mutual information maximization, encoding user preference and group representations as views for contrastive learning. Meanwhile, HHGR (Zhang et al., 2021a) constructs fine- and coarse-grained user-level hypergraphs, encoding augmented views of user nodes at different granularities. Moreover, CubeRec (Chen et al., 2022d) transforms the recommendation scenario using a hypercube framework, encoding user and item representations with LightGCN and treating each user representation as a view for pairing and learning. Lastly, SGGCF (Li et al., 2023d) adopts a data-based view creation approach, constructing a group-user-item graph, augmenting it with node/edge dropout, and encoding diverse views for each node.
  • Pair Sampling. In BunRec, users are linked with user groups, forming positive pairs between user and group representations. GroupIM maximizes mutual information (MI) between these representations. In HHGR and SGGCF, each graph node has multiple augmented views, which are naturally treated as positive pairs, while views of other nodes are considered negative samples. CubeRec constructs a group hypercube based on user representations, identifying the intersection between two group hypercubes. Each overlapping user is paired with the hypercube as a positive pair, and other users are considered negative samples. The distance between each user and the hypercube is then calculated for contrastive optimization.
  • Contrastive Objective. GroupIM utilizes a noise-contrastive objective with a binary cross-entropy loss, specifically a JS-based objective which maximizes the mutual information (MI) between positive pairs. On a related note, HHGR also employs a similar JS-based optimization approach. In CubeRec, an explicit margin loss is utilized to ensure that the distance-to-hypercube of positive pairs is smaller than that of negative pairs. For SGGCF, it adopts the InfoNCE-based objective to achieve the contrastive learning process.
@@ 4.8. Multi-behavior Recommendation
IICL (Li et al., 2023b) 2023 DASFAA Model-based Natural & Score-based InfoNCE-based
@@
TMCL (Xuan and Li, 2023) 2023 DASFAA Model-based Natural InfoNCE-based
@@
KMCLR (Xuan et al., 2023) 2023 WSDM Data & Model-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
VCGAE (Liu et al., 2023a) 2023 ICDM Variational Autoencoding Target Behavior Graph
@@
Liu et al. (Liu et al., 2022b) 2022 ICMR Model-based Natural JS-based
@@
MICRO (Zhang et al., 2022d) 2023 TKDE Model-based Natural InfoNCE-based
@@
GL Method Venue Year Generative Learning Paradigm Generation Target
MVGAE (Yi and Chen, 2021) 2022 TMM Variational Autoencoding Edge in Graph
@@ 4.8.1. Task Formulation
In Multi-behavior Recommendation (MbeRec), the user-item interaction is extended to incorporate behavior heterogeneity, resulting in a 3D tensor denoted as $\mathcal{A}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|\times|\mathcal{B}|}$, where =$\mathcal{B}=\{b_{1},...,b_{K}\}$ depicts the set of different types of behaviors (e.g., page view, add-to-cart). The MbeRec aims to provide item recommendations for the target behavior (e.g., purchase) by leveraging the diverse behavior information.
@@ 4.8.2. Contrastive Learning in MbeRec
In MbeRec, contrastive learning leverages multi-behavior user-item interactions to create diverse behavior-specific views for each data sample. These views, which differ among various methods, are ideal for contrastive learning. The methods are listed in Table 8.
@@
  • View Creation. Various methods employ model-based view creation and data-level view augmentation. HMG-CR (Yang et al., 2021a) uses hyper meta-paths to construct multiple hyper meta-graphs, encoding diverse behavior embeddings as views for each user node using distinct graph encoders. S-MBRec (Gu et al., 2022) partitions the multi-behavior graph into sub-graphs, encoding multiple GCN embeddings for each user/item as views. MMCLR (Wu et al., 2022b) encodes sequence and graph views for each user under different behaviors, obtaining a fusion view by combining these representations. CML (Wei et al., 2022) and IICL (Li et al., 2023b) leverage a behavior-aware graph neural network to encode user embeddings for each behavior, treating each behavior type as a separate view. MixMBR (Qiao et al., 2023) encodes behavior-specific embeddings and views through behavior message passing and model-parameter mixup. TMCL (Xuan and Li, 2023) utilizes temporal information to encode diverse behavior embeddings as distinct views through temporal-aware global learning. RCL (Wei et al., 2023b) encodes behavior-specific short-term and long-term interest embeddings as separate views, combining them to create a multi-behavior pattern view. MBSSL (Xu et al., 2023) employs behavior-aware GNNs to encode views for inter-behavior contrastive learning, and utilizes edge dropout to obtain augmented views for intra-behavior contrastive learning. KMCLR (Xuan et al., 2023) encodes behavior-specific views and employs knowledge-based data augmentation to obtain KG-aware views for each item for pairing.
@@
  • Pair Sampling. In HMG-CR, user behavior embeddings form positive pairs, with embeddings from other graphs being negative pairs. S-MBRec calculates similarity scores based on interaction nodes to determine positive user/item pairs. MMCLR, CML, MixMBR, TMCL, MBSSL and KMCLR treat different behavior views of the same node as positive pairs, and others as negative samples. In MBSSL, user similarity scores are computed to detect false negative samples and filter nagative pairs. Moreover, KMCLR also pairs two knowledge-augmented views of the same item for knowledge-aware contrastive learning. IICL uses intra- and inter-behavior contrastive learning, where even-layered embeddings are positive pairs in intra-behavior CL, and different behaviors for the same user are positive pairs in inter-behavior CL. Lastly, in RCL, the fused multi-behavior view is paired with other behavior-specific views as positive pairs for the same user, while considering other users as negative samples.
@@
  • Contrastive Objective. Most of the methods (Yang et al., 2021a; Gu et al., 2022; Wei et al., 2022; Li et al., 2023b; Qiao et al., 2023; Xuan and Li, 2023; Wei et al., 2023b; Xuan et al., 2023) in MbeRec utilize the InfoNCE objective to optimize contrastive learning. In MMCLR, it explicitly pulls the positive pairs closer together while pushing away negative pairs by ensuring that the similarity score of the former is larger than that of the latter.
4.8.3. Generative Learning in MbeRec

### v3

@@ 4.5.3. Reconstructive Learning in CroRec
  • Reconstructive Learning Paradigm in CroRec begins by employing deep generative latent variable models (i.e., variational auto-encoders) to encode the latent space of user preferences. Subsequently, it transitions to the denoising diffusion model to facilitate knowledge transfer across domains. Salah et al. propose RA/SA-VAE (Salah et al., 2021) to simultaneously fit the target observations and align the hidden space encoded by VAE with the source latent space. They introduce both rigid alignment and soft alignment techniques with varying degrees of user preference alignment. Furthermore, VDEA (Liu et al., 2022a) extends the alignment of latent spaces using variational autoencoding at both local and global levels, allowing for the exploitation of domain-invariant features across different domains, including both overlapped and non-overlapped users. Recently, DiffCDR (Xuan, 2024) leveraged denoising probabilistic models (DPMs) to process noisy data and generate denoised results, effectively transferring data between domains.
  • Reconstruction Target in cross-domain recommendation systems serves as a fundamental mechanism to bridge information gaps between domains by reconstructing missing or incomplete data representations. RA/SA-VAE and VDEA both adopt the user rating reconstruction paradigm to optimize the variational model. In these models, the reconstruction target is the user ratings for items, represented as vectors in the model. The obtained ratings include ratings for uninteracted items, which are subsequently used for ranking and recommendation. On the other hand, DiffCDR focuses on reconstructing user features in the target domain by reversing the diffusion process conditioned on the corresponding user’s information in the source domain.
@@ 4.5.4. Adversarial Learning in CroRec
  • Adversarial Learning Paradigm in CroRec leverages discriminative mechanisms to facilitate effective domain knowledge transfer. Several methods employ adversarial learning for cross-domain recommendation. For instance, Su et al. (Su et al., 2022) use a discriminator to identify domain-specific features, with gradients propagating to optimize the feature generator. Similarly, RecGURU (Li et al., 2022b) feeds encoded user representations from both domains into a discriminator for classification, enabling differentiable adversarial learning. Moreover, ACLCDR (Hsu et al., 2023) utilizes Deep Q-Learning to design a generator that enhances the interaction matrix with fake interactions, optimizing it using reward signals from downstream tasks. In addition, DA-CDR (Zhang et al., 2021b) encodes domain features of users and items to fool the discriminator, ensuring effective domain knowledge transfer through adversarial learning using a gradient reversal layer (GRL) for differentiable training. Furthermore, DA-DAN (Guo and other, 2023) incorporates adversarial domain adaptation into unsupervised non-overlapping CroRec, with the discriminator classifying the domain of encoded user features for overall differentiable training.
  • Discrimination Target in CroRec focuses on distinguishing domain-specific features to promote domain-invariant representation learning.. Su et al. (Su et al., 2022)’s domain adaptation method uses features from source and target domains to discriminate between them, forcing the generator to encode domain-invariant user preferences, thereby confusing the discriminator. RecGURU adopts a similar approach, discriminating encoded user representations from both domains to generalize user representation learning across domains. In ACLCDR, the augmented interaction matrix from DQN is used for inference in subsequent models, with the recommendation model acting as an implicit discriminator. DA-CDR employs a discriminator to classify domain features of users/items, encouraging the generator to encode additional domain-shared information for cross-domain recommendation. Lastly, DA-DAN encodes user-specific item sequence representations as domain features for the discriminator’s classification task, enabling the learning of domain-invariant features for recommendation.
4.5.5. In-depth Analysis of SSL-based Cross-domain Recommendation
In cross-domain recommendation, three main SSL approaches-contrastive learning, generative learning, and adversarial learning-each present unique strengths. Contrastive learning is effective in generating multiple views (Wang et al., 2021a; Zhang et al., 2022d) and enhancing mutual information across different domains, facilitating smooth knowledge transfer. Reconstructive learning, utilizing frameworks like variational autoencoders, aids in aligning latent spaces and reconstructing user preferences, offering a comprehensive representation of interactions (Salah et al., 2021; Liu et al., 2022a). Yet, it may face challenges in scalability with high-dimensional datasets. Adversarial learning focuses on producing domain-invariant features (Su et al., 2022; Li et al., 2022b) that improve generalization across domains, but its training can be unstable due to the competitive nature of the adversarial process involved.
@@ 4.6. Bundle Recommendation
RL Method Venue Year Reconstructive Learning Paradigm Reconstruction Target
DGMAE (Ren et al., 2023b) 2023 SIGIR Maksed Autoencoding Edge in User-Item Graph
@@
CubeRec (Chen et al., 2022b) 2022 SIGIR Model-based Score-based Explicit
@@ 4.6.2. Contrastive Learning in BunRec
In bundle recommendation with CL (Table 7), the intricate relationships between items and bundles, along with user-item interactions, are leveraged to construct user-bundle relationships. Various methodologies are subsequently designed to create contrastive views based on these multi-level relationships.
  • View Creation in BunRec contrastive learning methods focuses on constructing meaningful representations from multi-granular bundle-item relationships and user interaction patterns. MIDGN (Zhao et al., 2022) employs neural models to encode distinct representations for each user and bundle, treating them as separate views, while incorporating a graph disentangle module to encode intent-aware representations and utilizing a user-bundle graph with LightGCN for cross-view representation learning. CrossCBR (Ma et al., 2022) constructs the user-bundle graph by leveraging user-item and bundle-item relationships, subsequently encoding user and bundle representations through graph augmentation and embedding augmentation techniques in both bundle and item views. Consequently, each user and bundle node possesses dual views for contrastive learning, following a similar paradigm to MIDGN.
  • Pair Sampling in BunRec contrastive methods leverages the dual-view architecture to construct positive and negative sample pairs for effective contrast learning. Both MIDGN and CrossCBR encode two distinct views for each user and bundle in the dataset, where the dual views of a specific user/bundle node naturally constitute the positive pair, while views from different nodes serve as negative samples for contrastive comparison.
  • Contrastive Objective in BunRec systems aims to optimize multi-granular representations through self-supervised learning signals derived from bundle-item hierarchies. Both MIDGN and CrossCBR utilize the InfoNCE-based objective to optimize their models through self-supervised learning mechanisms.
4.6.3. Reconstructive Learning in BunRec
In BunRec, generative learning (Table 7) adopts auto-encoder paradigm to learn meaningful representations. The primary approach (i.e., DGMAE (Ren et al., 2023b)) is illustrated as follows.
  • Reconstructive Learning Paradigm leverages knowledge distillation combined with masked autoencoding mechanisms. In DGMAE, a teacher GNN model is first trained using user-bundle interactions, with its knowledge subsequently distilled into a student graph-masked auto-encoder model following mask autoencoding principles.
  • Reconstruction Target here focuses on masking and reconstructing critical structures to enhance model robustness. DGMAE employs an masking strategy on edges within the graph, prioritizing edges with lower sparsity scores for masking to increase reconstruction challenge and improve model robustness.
4.6.4. In-depth Analysis of SSL-based Bundle Recommendation
SSL techniques in this scenario address data sparsity and complex item-bundle synergies through distinct paradigms. CL-based methods (Zhao et al., 2022; Ma et al., 2022) enhance representation alignment across user-bundle and item-bundle views, improving robustness to data sparsity through multi-perspective augmentation. However, their effectiveness depends on domain-informed view design to balance semantic relevance and computational efficiency. Reconstructive approaches (Ren et al., 2023b) systematically recover masked user-item interactions, prioritizing high-affinity patterns to alleviate sparsity. A integration of them in the future could holistically address unique demands, such as capturing transient user intents and evolving bundle-item synergies.
@@ 4.7.2. Contrastive Learning in GroRec
  • View Creation in GroRec leverages multi-granular group structures and user-item interactions to construct meaningful representation views. GroupIM (Sankar et al., 2020) pioneers the use of mutual information maximization, encoding user preference and group representations as views for contrastive learning. Meanwhile, HHGR (Zhang et al., 2021a) constructs fine- and coarse-grained user-level hypergraphs, encoding augmented views of user nodes at different granularities. Moreover, CubeRec (Chen et al., 2022b) transforms the recommendation scenario using a hypercube framework, encoding user and item representations with LightGCN and treating each user representation as a view for pairing and learning. Lastly, SGGCF (Li et al., 2023d) adopts a data-based view creation approach, constructing a group-user-item graph, augmenting it with node/edge dropout, and encoding diverse views for each node.
  • Pair Sampling in GroRec constructs positive and negative sample pairs from group-user relationships and multi-view representations.. In BunRec, users are linked with user groups, forming positive pairs between user and group representations. GroupIM maximizes mutual information (MI) between these representations. In HHGR and SGGCF, each graph node has multiple augmented views, which are naturally treated as positive pairs, while views of other nodes are considered negative samples. CubeRec constructs a group hypercube based on user representations, identifying the intersection between two group hypercubes. Each overlapping user is paired with the hypercube as a positive pair, and other users are considered negative samples. The distance between each user and the hypercube is then calculated for contrastive optimization.
  • Contrastive Objective employs various optimization strategies tailored to group-aware recommendation scenarios.. GroupIM utilizes a noise-contrastive objective with a binary cross-entropy loss, specifically a JS-based objective which maximizes the mutual information (MI) between positive pairs. On a related note, HHGR also employs a similar JS-based optimization approach. In CubeRec, an explicit margin loss is utilized to ensure that the distance-to-hypercube of positive pairs is smaller than that of negative pairs. For SGGCF, it adopts the InfoNCE-based objective to achieve the contrastive learning process.
4.7.3. In-depth Analysis of SSL-based Group Recommendation
Contrastive learning-based methods have emerged as the mainstream approach in group recommendation due to their ability to effectively leverage user-item and user-group relationships. These methods, such as GroupIM (Sankar et al., 2020), HHGR (Zhang et al., 2021a), and SGGCF (Li et al., 2023d), excel in capturing fine-grained interaction patterns through multi-view representations, enabling robust modeling of the complex dynamics inherent in groups. However, CL-based methods are inherently constrained by their significant computational demands, which present a trade-off in fully optimizing group recommendation with CL. It would be promising to explore other SSL techniques, such as adversarial or reconstructive approaches, that may offer complementary advantages.
@@ 4.8. Multi-behavior Recommendation
IICL (Li et al., 2023f) 2023 DASFAA Model-based Natural & Score-based InfoNCE-based
@@
TMCL (Xuan et al., 2023b) 2023 DASFAA Model-based Natural InfoNCE-based
@@
KMCLR (Xuan et al., 2023a) 2023 WSDM Data & Model-based Natural InfoNCE-based
MENTOR (Xu et al., 2025) 2025 AAAI Data & Feature-based Natural InfoNCE-based
RL Method Venue Year Reconstructive Learning Paradigm Reconstruction Target
VCGAE (Liu et al., 2023b) 2023 ICDM Variational Autoencoding Target Behavior Graph
@@
Liu et al. (Liu et al., 2022c) 2022 ICMR Model-based Natural JS-based
@@
MICRO (Zhang et al., 2022a) 2023 TKDE Model-based Natural InfoNCE-based
@@
RL Method Venue Year Reconstructive Learning Paradigm Reconstruction Target
MVGAE (Yi et al., 2021) 2022 TMM Variational Autoencoding Edge in Graph
@@ 4.8.1. Task Formulation
In Multi-behavior Recommendation (MbeRec), the user-item interaction is extended to incorporate behavior heterogeneity, resulting in a 3D tensor denoted as $\mathcal{A}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|\times|\mathcal{B}|}$, where $\mathcal{B}=\{b_{1},...,b_{K}\}$ depicts the set of different types of behaviors (e.g., page view, add-to-cart). The MbeRec aims to provide item recommendations for the target behavior (e.g., purchase) by leveraging the diverse behavior information.
@@ 4.8.2. Contrastive Learning in MbeRec
In MbeRec, CL leverages multi-behavior user-item interactions to create behavior-specific views for effective representation learning. The methods are summarized in Table 8.
@@
  • View Creation in MbeRec focuses on constructing behavior-aware representations through model-based and data-level augmentation strategies. HMG-CR (Yang et al., 2021a) uses hyper meta-paths to construct multiple hyper meta-graphs with distinct encoders for behavior embeddings. S-MBRec (Gu et al., 2022) partitions multi-behavior graphs into sub-graphs, encoding multiple GCN embeddings as views. MMCLR (Wu et al., 2022b) encodes sequence and graph views under different behaviors, obtaining fusion views by combining representations. CML (Wei et al., 2022) and IICL (Li et al., 2023f) leverage behavior-aware GNNs to encode user embeddings for each behavior type as separate views. MixMBR (Qiao et al., 2023), TMCL (Xuan et al., 2023b), and RCL (Wei et al., 2023b) encode behavior-specific embeddings through parameter mixup, temporal-aware learning, and short/long-term interest modeling respectively. MBSSL (Xu et al., 2023) employs behavior-aware GNNs with edge dropout for augmented views, while KMCLR (Xuan et al., 2023a) incorporates knowledge-based data augmentation.
@@
  • Pair Sampling in MbeRec methods constructs positive and negative pairs from behavior-specific representations and similarity-based mechanisms. Most methods treat different behavior views of the same node as positive pairs, with other nodes serving as negative samples. HMG-CR pairs user behavior embeddings as positives, S-MBRec calculates interaction-based similarity scores for pairing, and MBSSL filters false negatives using user similarity scores. IICL employs both intra-behavior (even-layered embeddings) and inter-behavior (different behaviors for same user) contrastive learning, while RCL pairs fused multi-behavior views with behavior-specific views.
@@
  • Contrastive Objective in MbeRec systems primarily employs InfoNCE-based optimization to maximize behavior-aware representation learning. Most methods (Yang et al., 2021a; Gu et al., 2022; Wei et al., 2022; Li et al., 2023f; Qiao et al., 2023; Xuan et al., 2023b; Wei et al., 2023b; Xuan et al., 2023a) utilize the InfoNCE objective for contrastive optimization. MMCLR explicitly pulls positive pairs closer while pushing negative pairs apart by ensuring higher similarity scores for positive pairs.
4.8.3. Reconstructive Learning in MbeRec